2) Show that the cost function for a firm with the constant returns Cobb-Douglas production function 

y=Az1αz21−α


 is given by 

C(p,y)=yp1αp21−αB


 , where B is a function of A and 

α


 only. Derive the conditional input demands


1
Expert's answer
2022-03-24T16:11:16-0400

"Y=AZ_1^\alpha Z_2^{1-\alpha}"

"X=b_0L^{b_1}K^{b_2}"

"C=wL+rK"

"C=f(x)"

Maximize "X=b_0L^{b_1}K^{b_2}"

Subject to "C=wL+rK."

"\frac{\delta \phi}{\delta L}=b_1\frac{X}{L}-\lambda w=0"

"\frac{\delta \phi}{\delta K}=b_2\frac{X}{K}-\lambda r=0"

"\frac{\delta \phi}{\delta \lambda}=(C-wL-rK)=0"

"b_1\frac{X}{L}=\lambda w"

"b_2\frac{X}{K}=\lambda r"

"\frac{b_1}{b_2}.\frac{K}{L}=\frac{w}{r}"

"K=\frac{w}{r}.\frac{b_2}{b_1}L"

"X=b_0[(\frac{w}{r})(\frac {b_2}{b_1})]^{b_2}L^{(b_1+b_2)}"

"L=(\frac{rb_1}{wb_2})^\frac{b_2}{b_1+b_2}(\frac{X}{b_0})^\frac{1}{b_1+b_2}"

"K=\frac{w}{r}.\frac{b_2}{b_1}.L"

"K=(\frac{wb_2}{rb_1})^\frac{1}{b_1+b_2}(\frac{X}{b_0})^\frac{1}{b_1+b_2}"

"C=(\frac{1}{b_0})^\frac{1}{b_1+b_2}[w(\frac{rb_1}{wb_2})^\frac{b_2}{b_1+b_2}+r(\frac {wb_2}{rb_1})^\frac{b_1}{b_1+b_2}].X^\frac{1}{b_1+b_2}"

"C(p,y)=yP_1^\alpha P_1^{1-\alpha}B"